# Reservoir-Computing Model for Mapping and Forecasting Neuronal Interactions from Electrophysiological Data

### 23 Jan 2024 v2

@@ 2.3 Prediction of spatio-temporal response of the network
The connectivity analysis discussed in the previous subsection provides mapping of the network intrinsic connections. Here we evaluate the performance of the model to predict the response of the network, given a specific input (stimulation).
Considering that the model has undergone training based on the inherent spontaneous activity within the network, capturing the fundamental signal propagation between populations in response to minor perturbations. Upon acquisition of a trained model (Eq. (2.1), (4)), from Eq. (2.1), we posses the following operator (synaptic function):
$\tilde{\mathbf{y}}[n+1]=\hat{\mathcal{F}}\Bigl{\{}\tilde{\mathbf{y}}[n]\Bigr{\}}$ (8)
Figure 4: Illustration of the stimulus test designed to assess the RC model’s network response prediction performance to a local stimulus. On the left, an electrode map highlights the stimulated electrode (in red) and the responsive electrodes (in yellow), demonstrating spatial prediction. This is further evaluated against the experimental/simulated response using the ROC AUC metric. On the right, a detailed temporal response from all involved electrodes showcases spatio-temporal prediction. This is quantitatively compared to the experimental/simulation test using the $\bar{R}$ metric (refer to Eq. (9)).
Figure 5: An example of a response test to a specific stimulation protocol (NEST simulation data). A stimulus was given around a neuronal population found on the third row and third column of the map (designated with a yellow star). (a) The ground truth response (simulated in NEST), Predicted response and the error between them. Two types of comparisons are shown- the prediction in terms of amplitude, for which the channel-wise error is given by $\varepsilon$ (Eq. (28); and a binarized prediction to distinguish between responsive and unresponsive channels, which further was analyzed using the ROC curve AUC. For this example: $AUC=1$ at $\alpha=0.5$ and $m=50$. (b) The corresponding ROC curve. (c) a detailed map of each channel’s response in a fixed window of time (full time axis = $40ms$, time step (integration time) = $2ms$). The $y$ axis is the normalized ISR in arbitrary units. $\bar{R}=0.158$ at $\alpha=0.5$ and $m=50$.
Figure 6: An example of a response test to a specific stimulation protocol (MEA experiment data). (a), (b) and (c) share the same description of Fig. 5. The X’s correspond to channels that were discarded from analysis due to low activity. $\bar{R}=0.047$; $AUC=0.87$$\alpha=0.2$ and $m=50$.
@@
Hence, given an initial state $\tilde{\mathbf{u}}[0]=\mathbf{u}_{in}$, representing locally stimulated population, we obtain the predicted response of the network $\tilde{\mathbf{U}}$, by propagating the initial state for a number of time steps. For evaluating the prediction, we need to estimate the error between the predicted response $\tilde{\mathbf{U}}$ and the observed (ground truth) response $\mathbf{U}$. For this evaluation, we prepared a specific test data $\mathbf{U}$ for every tested neuronal network (experimental and simulated), corresponding to a localized stimulus applied in the vicinity of a specific neuronal circuit (in the experimental context- around a specific electrode).
In this context, we conducted a comprehensive evaluation of prediction quality through two distinct approaches: 1. ROC Curve Analysis: This method enabled us to spatially discriminate between responsive and unresponsive circuits. 2. Error evaluation between the full temporal profile of $\mathbf{U}$ and $\tilde{\mathbf{U}}$: To assess the predictive performance of temporal responses within network circuits, we introduced the following error function, based on the root-mean-squared-error (RMSE):
$\bar{R}=\sum_{i}\chi_{i}\min_{\tau}\left(\sqrt{\sum_{j}\chi_{j}\bigl{(}U_{i,j}-\tilde{U}_{i,j-\tau}\bigr{)}^{2}}\right)$ (9)
@@
Where $\chi_{i,j}$ are the spatial and temporal weighting coefficients respectively; $\tau\in[\tau_{min},\tau_{max}]$ is the time lag between the predicted and the observed responses. This metric quantified the accuracy of our predictions in capturing both the amplitude and time lag aspects of the actual responses, providing a comprehensive evaluation of temporal prediction quality. Fig. 4 illustrates the concept of only spatial and spatio-temporal predictions. Figs. 5 and 6 show an example of the response prediction test of the RC model using electrophysiological data from NEST simulation and MEA experiment, respectively. Figs. 7 and 8 show the summary of the results for the response prediction test in terms of $\bar{R}$ and AUC metrics for electrophysiological data from NEST simulation and MEA experiment, respectively.
Figure 7: Performance of the RC model for the network response prediction (NEST simulation data). The RC model underwent training using the basal activity of a specific network. Subsequently, the trained model was subjected to testing through local stimulation, and its response was compared against the corresponding ground truth response, as simulated in NEST. (a), (c) and (e) present the $\bar{R}$ metric (spatio-temporal response)- lower values are indicative of a more accurate prediction; (b), (d) and (f)- AUC metric (spatial response)- values closer to 1 indicate more accurate prediction. (a), (b) the metrics as a function of $q$ (Eq. (7)) the horizontal dashed line indicates the median value of the associated metric, the vertical line is at $q=100$, which we set as an arbitrary boundary between sparse and dense data. Each dot corresponds to the mean metric per experiment of a specific network (averaged over different stimulation protocols). It is noteworthy that denser datasets, characterized by large values of $q$, generally yield more accurate prediction outcomes. (c), (d) box plots of the metrics categorized by the simulated network size. (e), (f) the response test results plotted against the connectivity results of the corresponding network. A noticeable correlation between the two is observed ($\bar{R}$ vs. $\rho$: $R^{2}=0.060$; Response AUC vs CM AUC: $R^{2}=0.176$)
Figure 8: Performance of the RC model for the network response prediction (MEA experiments data). (a) and (b) share the same description of Figs. 6(a), 6(b).
Table 1: Summary of the response prediction test by the RC model. The table shows the results concerning the 2 metrics used: $\bar{R}$ (spatio-temporal prediction) and AUC (spatial prediction). The datasets originate from both NEST simulations and MEA experiments.
@@ 3 Discussion
In this work we developed a computational model which decodes spatio-temporal data from electro-physiological measurements of neuronal cultures, reconstructs the network structure on a macroscopic domain and can predict the response of a given network. The main aim of the model is to create an advanced experimental data analysis tool for processing complex time-series. The results obtained indicate that the model functions not just as a data analyzer but could also be used as a network simulator. Below we review the fundamentals of the model, we also attempt to give interpretation to the computational processes and we openly address some of the limitations we have encountered during our research.
The main idea on which the model has been developed relies primarily on the neural rate-coding of information paradigm [15], which states that firing rate (of action potentials) is the main carrier of information in the communication between the neuronal cells. The second presumption that establishes the model regards the essence of electrophysiological measurement of large neuronal populations, in particular extracellular measurements, which have been used to validate the model. As for the computational structure, the model is founded on a recurrent neural network (RNN) with RCN architecture. The rationale to use this kind of neural network for the current motives is rather self-evident. It has been shown in various studies that biological neuronal networks (e.g., the brain cortex) possess complex computational capacities, in particular temporal data is integrated in a recurrent manner and induce state-dependent synaptic transmission [16, 17, 18, 19, 20]. In other words, the network memorizes the previous impulses and modifies its synapses accordingly. The time-scale of this process varies according to the functionality of the network. RCN is a bio-inspired approach, where the neural recurrence and nonlinearity are taken into account. Another great advantage of this approach is the simplicity of the training, where only the output layer is trainable. The output layer can in principle be fully linear (but not necessarily), as the nonlinearity is embedded in the reservoir neurons. Utilizing a linear output layer gives the possibility to train the network with linear regression, which makes the process significantly cheaper (in terms of computational power) than training multilayered recurrent neural networks, such as Long-Short-Term-Memory (LSTM).
@@
RCN has been used in studies of electrophysiological signals from neuronal networks, in particular in these works the guiding principle is the exploitation of the reservoir-like properties of neuronal networks to adapt it to perform input-output tasks [8, 21, 7, 22, 23]. Our methodology, on the other hand, is to simulate the neuronal network using RCN approach to extract the network properties of a given culture such as the connectivity map of the network and its functionality on a macroscopic scale. Similar attitude towards this question was reported in [24] where the RCN approach was applied to learn the dynamics of spike-trains in neuronal cultures, in particular in MEA recordings. The work however focused on the prediction of spike-trains as a point-process propagation between distinctive input and output nodes, using the log-likelihood optimization. Additionally in that work, the authors implemented different types of neural networks, including an adaptive reservoir with a nonlinear output layer, which in practice transcends the modeling complexity discussed in this paper.
@@
In this work, we implement RCN as rate-coded spiking neural network, which we train to adapt the time-domain nonlinear synaptic function. The data that is used to train the model is rate-encoded spike trains. This data structure follows an assumption that the spikes that propagate within the network carry information in a characteristic time constant window (unique to each network) and is defined as the spike-integration time. In addition, we extract from the data, temporal events that exhibit significant spiking activity in short period of times (network bursts). As a result the time-series data undergoes a significant dimensionality reduction.
@@
The unique RCN design we have implemented decompose the information processing of the network at each time step (lasting one integration time in the real-world) to a pre-synaptic stage, executed by the input layer and the input of the reservoir layer; and the post-synaptic stage executed by the output layer and the output of the reservoir layer (Figs. 0(c),0(d)). This stage sequence correspond to one time-step and does not necessarily describe the realistic temporal propagation of signals within this time window. The reservoir layer consists of a number of independent micro-reservoirs which is equal to the number of the macroscopic network nodes (corresponding to the number of measurement sites or electrodes). In the pre-synaptic stage we sample the network state, each node independently by its own micro-reservoir. Each micro-reservoir integrates its current state with the input signal and applies nonlinear function. In the post-synaptic stage, the integrated and processed state of the reservoir (which is composed of all micro-reservoir states), is linearly mapped to the next network state, taking in consideration the synaptic coupling between the different micro-circuits.
The model, to some extent, is founded on the nonlinear leaky integrate-and-fire approach [5]. The structural design of our Reservoir Computing Network (RCN) adheres to distinct principles governing energy dynamics within the network. At its core, the structure is built upon the fundamental principles of reciprocity and energy conservation, augmented by the incorporation of a leaky memory and nonlinear elements. The dynamic behavior is intricately described through the interplay of information integration and synaptic energy exchange.
@@
We showed that by adapting the model to produce the dynamics observed in the experiments followed by applying a linear approximation of the transfer function, we obtain the fundamental (or intrinsic) connections between the populations of the network. These intrinsic connections describe the fundamental connectivity of the given network without internal or external stimuli that may modify them. Indeed, as seen from the equations describing the dynamics, during the full operation of the model these connections change depending on the activity of the network. The connectivity map we obtain from this analysis can be referred as Effective Connectivity rather than Functional Connectivity [25], because the resulting connections are derived from the actual causality relations between the populations based on the electrical signals; in contrast to the statistical relations seen in Cross-Correlation or Transfer Entropy analysis, which yield connectivity referred as Functional Connectivity and hence we assess the better performance of this model in predicting the connectivity of the network.
@@
Further, we tested the predictive capabilities of the model in terms of input-output prediction. Here we trained the model on the network basal (spontaneous) activity, with the assumption that this dataset would unveil the interactions among network nodes. During the testing phase, we posited that applying local stimulation to the network would elicit a correlated response from the circuits connected to the stimulated population. To assess this hypothesis, we gathered distinct datasets: one capturing spontaneous activity and another capturing evoked activity resulting from a local stimulus. The former was employed for training and validation, while the latter served as the test dataset. We subsequently simulated the model’s response to the specific stimulus and assessed the accuracy of our predictions. This procedure was carried out using both synthetic and experimental data. With ROC curve metrics, we have found that the model can spatially predict with high accuracy the network nodes that respond to the stimulus. However we also tested how accurate the prediction is in the temporal window of each node (electrode). Here we have found that the temporal profile of the predicted response may be distorted or time shifted. This could be due to the fact that some physical attributes of the stimulus are not taken into consideration of this model, such as the stimulus pulse shape (in amplitude and time) and hence the temporal dynamics is limited to time scales of the discretized time-step (defined as the integration time), nevertheless the temporal prediction in some cases reached high performance. The limitations in achieving precise temporal predictions may not solely be attributed to the model’s constraints, but also to the chosen training approach. For instance, the simplifying assumption that predicting evoked activity can be accomplished solely by learning from basal activity might prove to be overly simplistic. Nevertheless, the model is not confined to the specific training and testing approach involving spontaneous and evoked activities. In practical applications, any type of electrophysiological signal data, following the same data collection principles, can be employed to train the model and potentially improve its prediction abilities. For example, utilizing the evoked activity dataset can reveal the connections within the circuitry responsible for responding to the stimulus.
@@
As one would expect, the performance of the model is directly dependent on the data it is trained on. An important (but may not be sufficient) feature in the data is its spiking modality. The spiking data is expected to be modulated in such way that there should be a typical or a range of time that the network exhibits fast spiking events at numerous channels. These events are often seen as synchronous or quasi-synchronous [26, 27]. In this way, the data can be easily separated to temporal windows of these time events, while the silent phases are discarded. Each of these windows serve as data batch. This quality can be seen in the inter-spike interval (ISI) histogram, where a bimodal (or multi-modal) shape presents. Another element in the data, which affects the quality of the model performance is the presence of the causality between time-steps, discretized by a typical short-time window (time-bin), which we infer as the spike integration time of each circuit. The model in fact is trained to find this causality between the time-steps, and as the integration time is, in principle, a free parameter, we have found that the performance of the model is strongly dependent on this value. This implies that the network has a characteristic time constant (or quasi-constant) that defines its information processing time. We assume that this integration time, can be obtained from some attributes of the data. We have found that the most significant peak of inter-burst-interval (IBI), can serve as a good approximate of this value (see supplementary materials). This feature is also related to the stochasticity of the data, since it can determine the ability of the model to learn the spatio-temporal patterns. The characterization of the data degree of stochasticity is beyond the scope of this paper.
@@
Another limitation which we discuss here is the fact that biological systems can exhibit modifications in their structure in rather short times. Since the data analysed by the discussed model describe the setting of the tested culture within a limited time frame, it is possible that some information about the tested culture can be missed. Therefore for using this model for systems that may vary fastly in time, it is possible to apply the training for short time sessions, and eventually to see the development of the system.
In summary, we have developed a computational model which is based on RCN. The model decodes the spatio-temporal patterns of spike-data and reconstructs the connectivity map of the tested network. The model is also used to predict the response of the given network to local stimuli. In this paper we developed and tested the model on data from in-vitro neuronal electrophysiological signals recorded on 60 electrodes MEA. In-vitro studies of neurons give a simplified representation of the structure and functionality of these networks in living organisms [28, 29]. Such approach assists in decomposing the extremely complex structure of living brain into smaller functional blocks. We assess that the methodology developed in this work can be also applicable on data with higher spatial resolution (such as from HD-MEA [30]) and hence to give - up to interactions between single neurons. In addition the model is not restricted to analyze only neuronal signals, but can also be applied on different types of time traces.

### 05 Mar 2024 v3

@@ 3.2 Prediction of spatio-temporal response of the network
Here, we assess the performance of the RC model in estimating the response of a neuronal network when presented with specific inputs (stimulation). The RC model was trained based on the neuronal spontaneous activity. After training, we posses the operator $\hat{\mathcal{F}}$ (Eq. (3)), which characterizes the network dynamics. Our hypothesis is that the trained RC model can also predict the network’s response to localized stimuli. To facilitate the discussion, we label the network state as $\tilde{\mathbf{u}}[n]\equiv\tilde{\mathbf{y}}[n]$. Given an initial state $\tilde{\mathbf{u}}[0]=\mathbf{u}_{in}$, representing a locally stimulated neuronal population, we compute the predicted response of the network $\tilde{\mathbf{U}}\in\mathbb{R}^{N_{ch}\times N_{t}}$, by propagating the initial state for a number of time steps using the trained RC model operator $\hat{\mathcal{F}}$. To evaluate the prediction, we estimate the error between the predicted response $\tilde{\mathbf{U}}$ and the observed response $\mathbf{U}$, obtained either from the ground-truth network or through experiments. For this evaluation, we prepared a specific test data $\mathbf{U}$ for each tested neuronal network (both experimental and simulated), corresponding to a localized stimulus applied in the vicinity of a specific neuronal population (in the experimental context- around a specific electrode).
The evaluation of the prediction accuracy is performed by a ROC curve analysis to spatially discriminate between responsive and unresponsive neuronal populations (nodes). We also evaluate prediction accuracy based on temporal state evolutions. Specifically, we define the error $\bar{R}$ as:
$\bar{R}=\sum_{i}\chi_{i}\min_{\tau}\left(\sqrt{\sum_{j}\upsilon_{i,j,\tau}\bigl{(}U_{i,j}-\tilde{U}_{i,j-\tau}\bigr{)}^{2}}\right)\;,$ (9)
@@
where $\chi_{i}$ and $\upsilon_{i,j,\tau}$ are the spatial and temporal weighting coefficients, respectively, (defined in Eqs. 25 and 27, respectively); $\tau\in[\tau_{min},\tau_{max}]$ is the time lag between the predicted and the observed responses. $\bar{R}$, described in details in the Section 5.5.2, quantifies the accuracy of the trained RC model in capturing both the amplitude and time lag aspects of the actual responses. Supplementary 2 illustrates the concept of only spatial and spatio-temporal predictions.
In Section 3.1, we benchmarked the model’s performance using synthetic data. In this section, we expand our evaluation to include not only synthetic data but also experimental data recorded from cultured mice cortical neurons via MEA. The NEST simulation data plays a crucial role in enhancing our datasets and enabling a comprehensive assessment of response prediction. The outcomes of the response test, derived from the utilization of NEST simulation data, are presented in Figs. 5 and 6. The former showcases specific results for a particular network, while the latter provides a summary of performance metrics. Similarly, Figs. 7 and 8 exhibit similar results obtained from MEA experimental data. Table 1 summarizes the evaluation of performance metrics.
Table 1: Performance metrics for the response prediction test. $\bar{R}$ represents the root-mean-squared-errors, assessing the accuracy of spatio-temporal prediction while AUC denotes the area-under-the-curve in ROC analysis, evaluating the accuracy of the spatial prediction. For each experiment, we employed either a simulated network or a biological culture, subjecting them to various stimulation protocols. The datasets originate from both NEST simulations and MEA experiments.
@@
3.2.1 Results for the NEST simulation training data
The comparison of the responses of the networks in Fig. 2 to a local stimulus is shown in Fig. 5. The stimulation protocol excites the neuronal population probed by the 11th channel (as referenced in Fig. 2). The ground-truth data for the test set was extracted from the NEST simulation, while the RC model-predicted data was generated by operating the trained RC model with an input vector $\mathbf{u}_{in}$ and propagating it for 20 time steps. Fig. 4(a) illustrates that only one of the three uncoupled groups yields a response for both models. Specifically, the spatial distribution of the response amplitudes (top panels in Fig. 4(a)) shows a proper distribution with varying amplitudes and temporal profiles (see Fig. Supplementary 3). These differences are quantified in the error maps, where each node displays the results of the error estimations (Eq. (28)). A simpler comparison can be performed by a binary assessment of predictions, which channels are active or resting (bottom panels in Fig. 4(a)). The corresponding ROC curve analysis, depicted in Fig. 4(b), yields $AUC=1$. More details on the analysis are provided in Section 5.
((a))
((b))
Figure 5: Results of a response test for a specific stimulation protocol on the same network as Fig. 2. The RC model was trained on NEST simulation data with a memory strength $\alpha=0.5$ and a reservoir dimension $m=50$. A stimulus was given on the channel (node) labelled by a yellow star in the maps. (a, top panels) The amplitude response map for the ground truth network (left), the amplitude response map predicted by the RC model (centre) and the error maps for the RC model predictions as given by $\varepsilon$ (Eq. (28) (right). The color maps refer to the amplitudes of the channel (node) response or to the error value. (a, bottom panels) Ground truth network’s channel activity map (left), RC model node activity prediction map (centre), and error map evaluated as activity prediction correctness (right). The color code is given in the inset. (b) The ROC curve.
Similar response prediction tests were conducted on all the networks discussed in Section 3.1 by employing various stimulation protocols for each network configuration. Figure 6 provides an overview of the results. Figures 5(a) and 5(b) show the results as a function of the parameter $q$ (Eq. (8)). Each data point represents the average value across different stimulation protocols applied to a specific network. The horizontal dashed lines illustrate the overall average, as presented in Table 1. Additionally, a vertical dashed line is arbitrarily placed at $q=100$ to distinguish between dense and sparse networks. The red zone in Figure 5(b) represents the region between a completely incorrect prediction ($AUC=0$) and a random prediction ($AUC=0.5$). Notably, the RC model consistently outperforms this “red zone.” Furthermore, there is a discernible trend suggesting that better performance is more likely for higher values of $q$. Considering the results as a function of the network size (Figures 5(c) and 5(d)), this suggests that better results are obtained with longer training periods. Finally, we examined the influence of the accuracy in the connectivity map retrieval on the response prediction test (Figures 5(e) and 5(f)). A correlation between a low $\bar{R}$ value for a large Pearson correlation coefficient ($\rho$) or between a high $AUC$ value of the response test and a high $AUC$ of the connectivity retrieval is clear.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 6: Performance of the RC model for the network response prediction test (using NEST simulation training data). (a), (c) and (e) present the $\bar{R}$ metric (spatio-temporal response)- lower values are indicative of a more accurate prediction; (b), (d) and (f)- AUC metric (spatial response)- values closer to 1 indicate more accurate prediction. (a), (b) the metrics as a function of $q$ (Eq. (8)). The horizontal gray dashed line indicates the mean value of the associated metric, the vertical line is at $q=100$, which we set as an arbitrary boundary between dense and sparse network. Each dot corresponds to the average over different protocols of a specific network. The red zone in (b) indicates the range where the predictions fall below the level of random predictions ($AUC<0.5$). (c), (d) box plots of the metrics for different simulated network size. (e), (f) the response test results against the connectivity retrieval results of the same network. The red lines refer to a least square fit ($\bar{R}$ vs. $\rho$: $R^{2}=0.060$; Response AUC vs CM AUC: $R^{2}=0.176$), the bands refer to the 95 % confidence interval.
3.2.2 Results for the MEA experiments training data
A neuronal culture was measured using a 60-electrode MEA. Utilizing the basal activity, an ICM was retrieved as shown in Fig. 3. Subsequently, a stimulation test was conducted by exciting the neuronal culture using a specific electrode or a localized spot illumination (see Supplementary Materials), and recording the resulting map with the MEA. Simultaneously, a predictive response test was performed using the trained RC model, where the same network node was excited as discussed in Section 3.2.1. A comparison between the experimental MEA data and the RC model predictions is presented in Figs. 7 and 8. Notably, the RC model accurately predicts the spatio-temporal response of the neuronal culture to a localized stimulus ($\bar{R}=0.047$). This is evident in the activity maps and the single-node temporal evolutions (single-channel recordings for the experiment) shown in Fig. 7. The ROC analysis yields a high $AUC$ value of 0.88 in this specific case. Furthermore, repeated experiments with different neuronal culture preparations and stimulation protocols were analyzed using the RC model, and prediction tests were performed. Overall, the RC model captures the essence of the neuronal culture interconnections, accurately predicting how localized excitation induces a network response. Quantitative assessment of the predictions using the $\bar{R}$ and $AUC$ performance metrics is shown in Fig. 8 and summarized in Table 1. It is observed that the RC model accuracy, as a function of the $q$ parameter, yields better results for an average network size and testing time. Overall, lower $\bar{R}$ and $AUC$ values with respect to the case of synthetic data are observed which indicates a proper description of the active channel temporal signals (as indicated by the $\bar{R}$ value) and a partial capture of the intricate network in the neuronal culture (as indicated by the $AUC$ value).
((a))
((b))
((c))
Figure 7: Response test to a specific stimulation protocol for the neuronal culture analyzed in Fig. 2. The RC model was trained on a 60 electrodes MEA recording data with a memory strength $\alpha=0.2$ and a reservoir dimension $m=50$. (a) Activity maps for the experiment (left), for the predicted response (center) and for the error (right). The same description as in Fig. 4(a) applies. (b) The ROC curve. (c) Each node (channel for experiment) temporal response to a stimulus given at the node (channel) 14 (labeled by a yellow star). The time scale ($x$ axis) is given in the legend. Data are sampled with a time step of 2 ms. The $y$ axis is the normalized ISR in arbitrary units, scaled identically. From these time profiles $\bar{R}=0.047$ and $\bar{\tau}_{l}=6.46ms$ were estimated.
((a))
((b))
Figure 8: Performance of the RC model for the response prediction test applied to neuronal cultures (MEA experiments data). (a) and (b) the $\bar{R}$ and $AUC$ metrics as a function of $q$ (Eq. (8)). The horizontal gray dashed line indicates the mean value of the associated metric, the vertical line is at $q=100$, which we set as an arbitrary boundary between dense and sparse neuronal cultures. The red zone in (b) indicates the range where the predictions fall below the level of random predictions ($AUC<0.5$).
@@ 4 Discussion
In this study, we developed a computational model that decodes spatio-temporal data from electrophysiological measurements of neuronal cultures. The model reconstructs the network structure on a macroscopic domain and predicts the response to a localized stimulus. Our primary goal was to create an advanced experimental data analysis tool for processing complex time-series. The results obtained indicate that the model not only serves as a data analyzer but can also function as a network simulator. In the following sections, we delve into the model’s fundamentals, provide insights into the computational processes, and openly address some of the limitations encountered during our research.
The model’s core concept is rooted in the neural rate-coding of information paradigm, as proposed in [20]. According to this paradigm, the firing rate of action potentials serves as the primary carrier of information in communication between neuronal cells. The second fundamental assumption underlying the model pertains to the validity of electrophysiological measurements of large neuronal populations, particularly extracellular measurements, to capture their network structure. These measurements have been instrumental in validating the model.
Regarding its computational structure, the model is built upon a recurrent neural network (RNN) with a reservoir computing network (RCN) architecture. The rationale for choosing this type of neural network is straightforward. Numerous studies have demonstrated that biological neuronal networks (such as the brain cortex) exhibit complex computational capabilities. Specifically, temporal data is integrated in a recurrent manner, leading to state-dependent synaptic transmission [21, 22, 23, 24, 25]. In other words, the network memorizes past impulses and dynamically adjusts its synapses accordingly. The timescale of this process varies based on the network’s functionality. RCN, inspired by biological systems, incorporates neural recurrence and nonlinearity. An additional advantage lies in its training simplicity: only the output layer requires training. This output layer can be fully linear (although not necessarily), as the nonlinearity is embedded within the reservoir neurons. Utilizing a linear output layer enables cost-effective training through linear regression, as compared to more computationally intensive multilayered recurrent neural networks.
@@
RCNs have been widely used for studying electrophysiological signals from neuronal networks [8, 26, 7, 27, 28]. In particular, these studies leverage the reservoir-like properties of neuronal networks to adapt them for input-output tasks. Our methodology, however, takes a different approach. We simulate the neuronal network using the RCN framework to extract network properties from a given culture. This includes analyzing the connectivity map of the network and understanding its macroscopic functionality. A similar perspective was explored in [29], where the RCN approach was applied to learn the dynamics of spike trains in neuronal cultures, specifically in recordings from MEAs. In that work, the focus was on predicting spike trains as point-process propagation between distinct input and output nodes, using log-likelihood optimization. Notably, the authors implemented various types of neural networks, including an adaptive reservoir with a nonlinear output layer, which introduces modeling complexity beyond what we discuss in our paper.
@@
In our research, the RCN is implemented as a rate-coded spiking neural networks, which we train to adapt to time-domain nonlinear synaptic functions. The training data consists of rate-encoded spike trains. This data structure operates under the assumption that spikes propagating within the network carry information within a characteristic time window—unique to each network—defined as the spike-integration time. Additionally, we extract temporal events from the data that exhibit significant spiking activity over short periods (referred to as network bursts). Consequently, the time-series data undergoes substantial dimensionality reduction. Our unique RCN design dissects the network’s information processing at each time step (equivalent to one integration time in the real world). This decomposition occurs into two stages:
  1. 1.
    Pre-Synaptic Stage: Executed by the input layer and the input of the reservoir layer.
  2. 2.
    Post-Synaptic Stage: Executed by the output layer and the output of the reservoir layer (see Figs. 0(c) and Figure Supplementary 1).
@@
It’s important to note that this stage sequence corresponds to one time-step and does not necessarily describe the realistic temporal propagation of signals within this time window. The reservoir layer comprises a set of independent micro-reservoirs, equal in number to the macroscopic network nodes (corresponding to the measurement sites or electrodes). In the pre-synaptic stage, we sample the network state independently for each node using its own micro-reservoir. Each micro-reservoir integrates its current state with the input signal and applies a nonlinear function. In the post-synaptic stage, the integrated and processed state of the reservoir (composed of all micro-reservoir states) is linearly mapped to the next network state, taking into account the synaptic coupling between different micro-circuits.
We demonstrated that by adapting the model to replicate the observed dynamics in our experiments, followed by applying a linear approximation of the transfer function, we can uncover the fundamental (or intrinsic) connections between the populations within the network. These intrinsic connections describe the network’s fundamental connectivity, independent of any internal or external stimuli that might modify them.
@@
Interestingly, as indicated by the equations governing the dynamics, these connections change during the full operation of the model, depending on the network’s activity. The resulting connectivity map from this analysis can be referred to as Effective Connectivity, in contrast to Functional Connectivity [30]. The term Effective Connectivity emphasizes that these connections are derived from actual causality relations between populations based on electrical signals, as opposed to the statistical relations seen in Cross-Correlation or Transfer Entropy analysis, which yield Functional Connectivity. We assessed the better performance of our model in predicting the connectivity of the network compared to these last methods.
@@
Furthermore, we tested the model’s predictive capabilities in terms of input-output prediction. We trained the model using network basal (spontaneous) activity, assuming that this dataset would reveal interactions among network nodes. During the testing phase, we hypothesized that applying local stimulation to the network would elicit a correlated response from the circuits connected to the stimulated population. To assess this hypothesis, we collected distinct datasets: one capturing spontaneous activity and another capturing evoked activity resulting from a local stimulus. The former was used for training and validation, while the latter served as the test dataset. Subsequently, we simulated the model’s response to the specific stimulus and evaluated the accuracy of our predictions. We conducted this procedure using both synthetic and experimental data. Our analysis, based on ROC curve metrics, revealed that the model can accurately predict the network nodes that respond to the stimulus in spatial terms. However, we also assessed the accuracy of temporal predictions for each node (electrode). We observed that the temporal profile of the predicted response may be distorted or time-shifted. This discrepancy could be attributed to the fact that certain physical attributes of the stimulus, such as pulse shape (amplitude and time), are not fully considered by the model. Consequently, the temporal dynamics are limited to the time scales of the discretized time-step (defined as the integration time). Nevertheless, in some cases, the temporal predictions achieved high performance. It’s important to note that the limitations in achieving precise temporal predictions may not solely stem from the model’s constraints but also from our chosen training approach. For example, the simplifying assumption that predicting evoked activity can be accomplished solely by learning from basal activity might prove overly simplistic. Nonetheless, the model is not confined to the specific training and testing approach involving spontaneous and evoked activities. In practical applications, any type of electrophysiological signal data, following the same data collection principles, can be employed to train the model and potentially improve its prediction abilities. For example, utilizing the evoked activity dataset can reveal the connections within the circuitry responsible for responding to the stimulus.
@@
The model’s performance directly depends on the data it is trained on. An essential feature in the data is its spiking modality, which should exhibit fast spiking events at specific time intervals. These events are often seen as synchronous or quasi-synchronous [31, 32]. By separating the data into temporal windows during these time events and discarding silent phases, each window serves as a data batch. This quality is evident in the inter-spike interval (ISI) histogram, which typically presents a bimodal or multi-modal shape. Another critical factor affecting model performance is the causality between time-steps, discretized by a short-time window (time-bin), which we infer as the spike integration time of each circuit. The model is trained to identify this causality, and the integration time significantly impacts its effectiveness. This implies that the network has a characteristic time constant defining its information processing time. We propose that this integration time can be approximated using attributes from the data, such as the most significant peak in the inter-burst interval (IBI). This feature also relates to the data’s stochasticity, influencing the model’s ability to learn spatio-temporal patterns. However, characterizing the degree of data stochasticity is beyond the scope of this paper.
@@
Another limitation of the model is that biological systems can undergo structural modifications relatively quickly. The data analyzed by the model describe the state of the tested culture within a limited time frame. Consequently, there is a possibility that some information about the tested culture might be overlooked. To address this, when using the model for systems that exhibit rapid variations over time, one can apply short training sessions and observe the system’s development.
Finally, our model was applied to in-vitro cultures of neurons, providing a simplified representation of the structure and functionality of these networks in living organisms [33, 34]. This approach aids in breaking down the intricate architecture of the living brain into smaller functional units. We believe that the methodology developed in this study can also be extended to data with higher spatial resolution, such as that obtained from HD-MEA [35], allowing for analysis of interactions even at the level of individual neurons. Furthermore, the model is not limited to analyzing only neuronal signals; it can also be applied to various other types of time traces.